# The equilibrium properties of obvious strategy profiles

### v1

@@ 5.3 Proof of Theorem 1
$\left|u_{i}^{n}(\mu,g^{n}_{-i})-u^{n}_{i}\Bigl{(}\mu,\lambda^{n}(i)\mu+\sum\limits_{j\in I^{n}\backslash\{i\}}\lambda^{n}(j)g^{n}(j)\Bigr{)}\right|\leq\frac{\varepsilon}{4}.$
Step 3. In this step, we estimate $\Bigl{|}u_{i}^{n}(g^{n})-u_{i}^{n}\bigl{(}g^{n}_{i},s(\widetilde{g})\bigr{)}\Bigr{|}$. Similarly, let $\{x^{n}_{i}\}_{i\in I^{n},n\in\mathbb{Z}_{+}}$ represent all the players’ strategies in games $\{G^{n}\}_{n\in\mathbb{Z}_{+}}$. The payoff of player $i\in I^{n}$ in game $G^{n}$ with the strategy profile $g^{n}$ can be rewritten as
$u_{i}^{n}(g^{n})=\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x^{n}_{i},\sum\limits_{j\in I^{n}}\lambda^{n}(j)\delta_{x^{n}_{j}}\bigr{)}\Bigr{]}.$
Since there exists some player $i^{\prime}\in I$ such that $G(i^{\prime})=G^{n}(i)$ (i.e., $u_{i^{\prime}}=u^{n}_{i}$), the payoff of player $i^{\prime}$ in large game $G$ with strategy profile $\widetilde{g}$ is equivalent to
$\int_{A}u_{i^{\prime}}\bigl{(}a,s(\widetilde{g})\bigr{)}\widetilde{g}(i^{\prime},\operatorname{d\!}a)=\int_{A}u_{i}^{n}\bigl{(}a,s(\widetilde{g})\bigr{)}g^{n}(i,\operatorname{d\!}a)=\mathbb{E}\bigl{[}u_{i}^{n}\bigl{(}x^{n}_{i},s(\widetilde{g})\bigr{)}\bigr{]}.$
@@
$\left|\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x^{n}_{i},\sum\limits_{j\in I^{n}}\lambda^{n}(j)\delta_{x^{n}_{j}}\bigr{)}\Bigr{]}-\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x^{n}_{i},s(\widetilde{g})\bigr{)}\Bigr{]}\right|$
$\leq\left|\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x^{n}_{i},\sum\limits_{j\in I^{n}}\lambda^{n}(j)\delta_{x^{n}_{j}}\bigr{)}\Bigr{]}-\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x^{n}_{i},\sum\limits_{j\in I}\lambda^{n}(j)g^{n}(j)\bigr{)}\Bigr{]}\right|$ (i)
$+\left|\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x^{n}_{i},\sum\limits_{j\in I^{n}}\lambda^{n}(j)g^{n}(j)\bigr{)}\Bigr{]}-\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x^{n}_{i},s(\widetilde{g})\bigr{)}\Bigr{]}\right|.$ (ii)
Estimation of (i). By step 2, for any $\varepsilon>0$, there exists a sequence of sets $\{S^{n}\}_{n\in\mathbb{Z}_{+}}$ such that $\lambda^{n}(S^{n})>1-\frac{\varepsilon}{2}$ for all $n\in\mathbb{Z}_{+}$, and a number $N_{\varepsilon}\in\mathbb{Z}_{+}$ such that for all $n\geq N_{\varepsilon}$, $i\in S^{n}$, we have $\Bigl{|}u_{i}^{n}(g^{n})-u_{i}^{n}\bigl{(}g^{n}(i),\sum_{j\in I^{n}}\lambda^{n}(j)g^{n}(j)\bigr{)}\Bigr{|}<\frac{\varepsilon}{4}$. That is,
$\left|\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x^{n}_{i},\sum\limits_{j\in I^{n}}\lambda^{n}(j)\delta_{x^{n}_{j}}\bigr{)}\Bigr{]}-\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x^{n}_{i},\sum\limits_{j\in I}\lambda^{n}(j)g^{n}(j)\bigr{)}\Bigr{]}\right|<\frac{\varepsilon}{4}.$
Estimation of (ii). Since all the functions in $\{u_{i}^{n}\}_{i\in S^{n},n\in\mathbb{Z}_{+}}$ are uniformly bounded by $M_{\varepsilon}$ and equicontinuous, for the given $\varepsilon>0$, there exists $\eta>0$ such that for any $\tau,\widetilde{\tau}\in\mathcal{M}(A)$ with $\rho(\tau,\widetilde{\tau})<\eta$, we have $|u(a,\tau)-u(a,\widetilde{\tau})|<\frac{\varepsilon}{4(2M_{\varepsilon}+1)}$, for all $a\in A,u\in\{u_{i}^{n}\}_{i\in S^{n},n\in\mathbb{Z}_{+}}$. For any bounded continuous function $h\colon A\to R$, let $\phi(u)=\int_{A}h(a)\overline{g}(u,\operatorname{d\!}a)$ for all $u\in\mathrm{supp}\lambda G^{-1}\subset\mathcal{U}_{A}$. Since $\overline{g}$ is continuous for $\lambda G^{-1}$-almost all $u\in\mathrm{supp}\lambda G^{-1}$, we know that $\phi$ is bounded and continuous for $\lambda G^{-1}$-almost all $u\in\mathrm{supp}\lambda G^{-1}$. Moreover, as $\lambda^{n}(G^{n})^{-1}$ converges weakly to $\lambda G^{-1}$ and $\mathrm{supp}\lambda^{n}(G^{n})^{-1}\subset\mathrm{supp}\lambda G^{-1}$, by Portmanteau Theorem (Klenke (2014, Theorem 13.16)), we have that
$\int_{\mathrm{supp}\lambda G^{-1}}\phi(u)\operatorname{d\!}\lambda^{n}(G^{n})^{-1}(u)\to\int_{\mathrm{supp}\lambda G^{-1}}\phi(u)\operatorname{d\!}\lambda G^{-1}(u).$
@@
$\int_{\mathrm{supp}\lambda G^{-1}}\phi(u)\operatorname{d\!}\lambda^{n}(G^{n})^{-1}(u)=\int_{I^{n}}\phi\bigl{(}G^{n}(i)\bigr{)}\operatorname{d\!}\lambda^{n}(i),$
@@
$\phi\bigl{(}G^{n}(i)\bigr{)}=\int_{A}h(a)\overline{g}\bigl{(}G^{n}(i),\operatorname{d\!}a\bigr{)}=\int_{A}h(a)g^{n}(i,\operatorname{d\!}a),$
@@
$\int_{I^{n}}\int_{A}h(a)g^{n}(i,\operatorname{d\!}a)\to\int_{I}\int_{A}h(a)\widetilde{g}(i,\operatorname{d\!}a),$
@@
by changing of variables. Therefore, $s(g^{n})$ converges weakly to $s(\widetilde{g})$, and there exists $\widetilde{N}_{1}\in\mathbb{Z}_{+}$ such that for all $n\geq\widetilde{N}_{1}$, $\rho\bigl{(}s(g^{n}),s(\widetilde{g})\bigr{)}<\eta$. Thus for all $n\geq\widetilde{N}_{1}$, $i\in S^{n}$, we have
$\left|\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x^{n}_{i},\sum\limits_{j\in I^{n}}\lambda^{n}(j)g^{n}(j)\bigr{)}\Bigr{]}-\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x^{n}_{i},s(\widetilde{g})\bigr{)}\Bigr{]}\right|\leq\dfrac{\varepsilon}{4(2M_{\varepsilon}+1)}.$
@@
$\left|\mathbb{E}\Bigl{[}u_{i}^{n}\Bigl{(}x^{n}_{i},\sum\limits_{j\in I^{n}}\lambda^{n}(j)\delta_{x^{n}_{j}}\Bigr{)}\Bigr{]}-\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x^{n}_{i},s(\widetilde{g})\bigr{)}\Bigr{]}\right|\leq\frac{\varepsilon}{4}+\dfrac{\varepsilon}{4(2M_{\varepsilon}+1)},$
for all $n\geq\max\{\widetilde{N}_{1},N_{\varepsilon}\}$, $i\in S^{n}$. Therefore, for all $n\geq\max\{\widetilde{N}_{1},N_{\varepsilon}\}$, $i\in S^{n}$, we have
@@
Step 4. We estimate $\Bigl{|}u_{i}^{n}(\mu,g_{-i}^{n})-u_{i}^{n}\bigl{(}\mu,s(\widetilde{g})\bigr{)}\Bigr{|}$ in this step. By using the triangle inequality, we have that
$\left|\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x_{\mu},\lambda^{n}(i)\delta_{x_{\mu}}+\sum\limits_{j\in I^{n}\backslash\{i\}}\lambda^{n}(j)\delta_{x^{n}_{i}}\bigr{)}\Bigr{]}-\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x_{\mu},s(\widetilde{g})\bigr{)}\Bigr{]}\right|$
$\leq\left|\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x_{\mu},\lambda^{n}(i)\delta_{x_{\mu}}+\sum\limits_{j\in I^{n}\backslash\{i\}}\lambda^{n}(j)\delta_{x^{n}_{j}}\bigr{)}\Bigr{]}-\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x_{\mu},\lambda^{n}(i)\mu+\sum\limits_{j\in I^{n}\backslash\{i\}}\lambda^{n}(j)g^{n}(j)\bigr{)}\Bigr{]}\right|$ (I)
$+\left|\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x_{\mu},\lambda^{n}(i)\mu+\sum\limits_{j\in I^{n}\backslash\{i\}}\lambda^{n}(j)g^{n}(j)\bigr{)}\Bigr{]}-\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x_{\mu},\sum\limits_{j\in I^{n}}\lambda^{n}(j)g^{n}(j)\bigr{)}\Bigr{]}\right|$ (II)
$+\left|\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x_{\mu},\sum\limits_{j\in I^{n}}\lambda^{n}(j)g^{n}(j)\bigr{)}\Bigr{]}-\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x_{\mu},s(\widetilde{g})\bigr{)}\Bigr{]}\right|$ (III)
for all $\mu\in\mathcal{M}(A)$, where $x_{\mu}$ is the random variable which induces the distribution $\mu$. The estimation of part (I) is the same as part (i) in step 3, while the estimation of part (III) is the same as part (ii) in step 3. Hence there exists $\widetilde{N}_{2}\in\mathbb{Z}_{+}$ such that for all $n\geq\widetilde{N}_{2}$, $i\in S^{n}$, $\mu\in\mathcal{M}(A)$,
@@
Below we estimate part (II). By the definition of the Prohorov metric $\rho$, we have that for any $\mu\in\mathcal{M}(A)$, $i\in I^{n}$,
$\rho\bigl{(}s(g^{n}),s(\mu,g^{n}_{-i})\bigr{)}\leq\sup\limits_{j\in I^{n}}\lambda^{n}(j).$
Since $\sup_{j\in I^{n}}\lambda^{n}(j)\rightarrow 0$, there exists $\widetilde{N}_{3}\in\mathbb{Z}_{+}$ such that for any $n\geq\widetilde{N}_{3}$, $\sup_{j\in I^{n}}\lambda^{n}(j)<\eta$. Thus, for any $n\geq\widetilde{N}_{3}$, $i\in I^{n}$, $\mu\in\mathcal{M}(A)$, $\rho\bigl{(}s(g^{n}),s(\mu,g^{n}_{-i})\bigr{)}<\eta.$ Recall that for all $i\in S^{n}$, $u_{i}^{n}$ is uniformly bounded by $M_{\varepsilon}$ and equicontinuous. As we proved in part (ii), we have that
$\left|\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x_{\mu},\lambda^{n}(i)\mu+\sum\limits_{j\in I^{n}\backslash\{i\}}\lambda^{n}(j)g^{n}(j)\bigr{)}\Bigr{]}-\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x_{\mu},\sum\limits_{j\in I^{n}}\lambda^{n}(j)g^{n}(j)\bigr{)}\Bigr{]}\right|\leq\dfrac{\varepsilon}{4(2M_{\varepsilon}+1)},$
for all $n\geq\widetilde{N}_{3}$, $i\in S^{n}$, $\mu\in\mathcal{M}(A)$. Thus for all $n\geq\max\{\widetilde{N}_{2},\widetilde{N}_{3}\}$, $i\in S^{n}$, $\mu\in\mathcal{M}(A)$,
$\left|\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x_{\mu},\lambda^{n}(i)\delta_{x_{\mu}}+\sum\limits_{j\in I^{n}\backslash\{i\}}\lambda^{n}(j)\delta_{x^{n}_{j}}\bigr{)}\Bigr{]}-\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x_{\mu},s(\widetilde{g})\bigr{)}\Bigr{]}\right|\leq\frac{\varepsilon}{4}+\dfrac{\varepsilon}{2(2M_{\varepsilon}+1)}.$
@@
Step 5. For any $n\geq\max\{N_{\varepsilon},\widetilde{N}_{1},\widetilde{N}_{2},\widetilde{N}_{3}\}$, $i\in S^{n}$, $\mu\in\mathcal{M}(A)$, we have
$\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x^{n}_{i},\sum\limits_{j\in I^{n}}\lambda^{n}(j)\delta_{x^{n}_{j}}\bigr{)}\Bigr{]}$
@@
$\geq\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x_{\mu},\lambda^{n}(i)\delta_{x_{\mu}}+\sum\limits_{j\in I^{n}\backslash\{i\}}\lambda^{n}(j)\delta_{x^{n}_{j}}\bigr{)}\Bigr{]}-\frac{\varepsilon}{2}-\frac{\varepsilon}{2(2M_{\varepsilon}+1)}.$
The first inequality follows from step 3 and the last inequality follows from step 4. The second inequality is due to the fact that $\widetilde{g}$ is a Nash equilibrium of $G$ such that for any player $i\in I$, her strategy $\widetilde{g}(i)$ is a best response with respect to the society summary $s(\widetilde{g})$, and $G^{n}(i)\in G(I)$ for all $i\in I^{n}$, $n\in\mathbb{Z}_{+}$. Hence we have that
$\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x^{n}_{i},\sum\limits_{j\in I^{n}}\lambda^{n}(j)\delta_{x^{n}_{j}}\bigr{)}\Bigr{]}\geq\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x_{\mu},\lambda^{n}(i)\delta_{x_{\mu}}+\sum\limits_{j\in I^{n}\backslash\{i\}}\lambda^{n}(j)\delta_{x^{n}_{j}}\bigr{)}\Bigr{]}-\varepsilon,$
for all $n\geq\max\{N_{\varepsilon},\widetilde{N}_{1},\widetilde{N}_{2},\widetilde{N}_{3}\}$, $i\in S^{n}$, $\mu\in\mathcal{M}(A)$. Thus we conclude that $g^{n}$ is an $\varepsilon$-Nash equilibrium of $G^{n}$, for all $n\geq\max\{N_{\varepsilon},\widetilde{N}_{1},\widetilde{N}_{2},\widetilde{N}_{3}\}$.
Given any sequence $\{\varepsilon_{n}\}_{n\in\mathbb{Z}_{+}}$ such that $\varepsilon_{n}>0$ for all $n\in\mathbb{Z}_{+}$ and $\{\varepsilon_{n}\}_{n\in\mathbb{Z}_{+}}$ converges to 00, there exists a strictly increasing sequence $\{\overline{N}_{n}\}_{n\in\mathbb{Z}_{+}}$ such that $\overline{N}_{n}\in\mathbb{Z}_{+}$ for all $n\in\mathbb{Z}_{+}$, and $g^{m}$ is a $\varepsilon_{n}$-Nash equilibrium of $G^{m}$ for all $m\geq\overline{N}_{n}$. Let $\varepsilon_{k}=\varepsilon_{n}$ if $\overline{N}_{n}\leq k<\overline{N}_{n+1}$, for all $n,k\in\mathbb{Z}_{+}$. Thus we have that $\{\varepsilon_{k}\}_{k\in\mathbb{Z}_{+}}$ converges to 00, and $g^{k}$ is an $\varepsilon_{k}$-Nash equilibrium of $G^{k}$ for all $k\geq\overline{N}_{1}$, which completes our proof of Theorem 1.

### v2

@@ 5.3 Proof of Theorem 1
[…]
$u_{i}^{n}(g^{n})=\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x^{n}_{i},\sum\limits_{j\in I_{n}}\lambda_{n}(j)\delta_{x^{n}_{j}}\bigr{)}\Bigr{]}.$
Since there exists some player $i^{\prime}\in I$ such that $G(i^{\prime})=G_{n}(i)$ (i.e., $u_{i^{\prime}}=u^{n}_{i}$), the payoff of player $i^{\prime}$ in large game $G$ with strategy profile $\widetilde{g}$ is equivalent to
[…]
@@
[…]
$\leq\left|\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x^{n}_{i},\sum\limits_{j\in I_{n}}\lambda_{n}(j)\delta_{x^{n}_{j}}\bigr{)}\Bigr{]}-\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x^{n}_{i},\sum\limits_{j\in I}\lambda_{n}(j)g^{n}(j)\bigr{)}\Bigr{]}\right|$ (i)
[…]
Estimation of (i). By step 2, for any $\varepsilon>0$, there exists a sequence of sets $\{S_{n}\}_{n\in\mathbb{Z}_{+}}$ such that $\lambda_{n}(S_{n})>1-\frac{\varepsilon}{2}$ for all $n\in\mathbb{Z}_{+}$, and a number $N_{\varepsilon}\in\mathbb{Z}_{+}$ such that for all $n\geq N_{\varepsilon}$, $i\in S_{n}$, we have $\Bigl{|}u_{i}^{n}(g^{n})-u_{i}^{n}\bigl{(}g^{n}(i),\sum_{j\in I_{n}}\lambda_{n}(j)g^{n}(j)\bigr{)}\Bigr{|}<\frac{\varepsilon}{4}$. That is,
$\left|\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x^{n}_{i},\sum\limits_{j\in I_{n}}\lambda_{n}(j)\delta_{x^{n}_{j}}\bigr{)}\Bigr{]}-\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x^{n}_{i},\sum\limits_{j\in I}\lambda_{n}(j)g^{n}(j)\bigr{)}\Bigr{]}\right|<\frac{\varepsilon}{4}.$
Estimation of (ii). Since all the functions in $\{u_{i}^{n}\}_{i\in S_{n},n\in\mathbb{Z}_{+}}$ are uniformly bounded by $M_{\varepsilon}$ and equicontinuous, for the given $\varepsilon>0$, there exists $\eta>0$ such that for any $\tau,\widetilde{\tau}\in\mathcal{M}(A)$ with $\rho(\tau,\widetilde{\tau})<\eta$, we have $|u(a,\tau)-u(a,\widetilde{\tau})|<\frac{\varepsilon}{4(2M_{\varepsilon}+1)}$, for all $a\in A,u\in\{u_{i}^{n}\}_{i\in S_{n},n\in\mathbb{Z}_{+}}$. For any bounded continuous function $h\colon A\to R$, let $\phi(u)=\int_{A}h(a)\overline{g}(u,\operatorname{d\!}a)$ for all $u\in\mathrm{supp}\lambda G^{-1}\subset\mathcal{U}_{A}$. Since $\overline{g}$ is continuous for $\lambda G^{-1}$-almost all $u\in\mathrm{supp}\lambda G^{-1}$, we know that $\phi$ is bounded and continuous for $\lambda G^{-1}$-almost all $u\in\mathrm{supp}\lambda G^{-1}$. Moreover, as $\lambda_{n}{G_{n}}^{-1}$ converges weakly to $\lambda G^{-1}$ and $\mathrm{supp}\lambda_{n}{G_{n}}^{-1}\subset\mathrm{supp}\lambda G^{-1}$, by Portmanteau Theorem (Klenke (2014, Theorem 13.16)), we have that
$\int_{\mathrm{supp}\lambda G^{-1}}\phi(u)\operatorname{d\!}\lambda_{n}{G_{n}}^{-1}(u)\to\int_{\mathrm{supp}\lambda G^{-1}}\phi(u)\operatorname{d\!}\lambda G^{-1}(u).$
@@
$\int_{\mathrm{supp}\lambda G^{-1}}\phi(u)\operatorname{d\!}\lambda_{n}{G_{n}}^{-1}(u)=\int_{I_{n}}\phi\bigl{(}G_{n}(i)\bigr{)}\operatorname{d\!}\lambda_{n}(i),$
@@
$\phi\bigl{(}G_{n}(i)\bigr{)}=\int_{A}h(a)\overline{g}\bigl{(}G_{n}(i),\operatorname{d\!}a\bigr{)}=\int_{A}h(a)g^{n}(i,\operatorname{d\!}a),$
@@
$\int_{I_{n}}\int_{A}h(a)g^{n}(i,\operatorname{d\!}a)\operatorname{d\!}\lambda_{n}(i)\to\int_{I}\int_{A}h(a)\widetilde{g}(i,\operatorname{d\!}a)\lambda(i),$
@@
by changing of variables. Therefore, $s(g^{n})$ converges weakly to s(g~)s($\widetilde${g}), and there exists $\widetilde{N}_{1}\in\mathbb{Z}_{+}$ such that for all $n\geq\widetilde{N}_{1}$, $\rho\bigl{(}s(g^{n}),s(\widetilde{g})\bigr{)}<\eta$. Thus for all $n\geq\widetilde{N}_{1}$, $i\in S_{n}$, we have
$\left|\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x^{n}_{i},\sum\limits_{j\in I_{n}}\lambda_{n}(j)g^{n}(j)\bigr{)}\Bigr{]}-\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x^{n}_{i},s(\widetilde{g})\bigr{)}\Bigr{]}\right|\leq\dfrac{\varepsilon}{4(2M_{\varepsilon}+1)}.$
@@
$\left|\mathbb{E}\Bigl{[}u_{i}^{n}\Bigl{(}x^{n}_{i},\sum\limits_{j\in I_{n}}\lambda_{n}(j)\delta_{x^{n}_{j}}\Bigr{)}\Bigr{]}-\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x^{n}_{i},s(\widetilde{g})\bigr{)}\Bigr{]}\right|\leq\frac{\varepsilon}{4}+\dfrac{\varepsilon}{4(2M_{\varepsilon}+1)},$
for all $n\geq\max\{\widetilde{N}_{1},N_{\varepsilon}\}$, $i\in S_{n}$. Therefore, for all $n\geq\max\{\widetilde{N}_{1},N_{\varepsilon}\}$, $i\in S_{n}$, we have
@@
[…]
for all $\mu\in\mathcal{M}(A)$, where $x_{\mu}$ is the random variable which induces the distribution $\mu$. The estimation of part (I) is the same as part (i) in step 3, while the estimation of part (III) is the same as part (ii) in step 3. Hence there exists $\widetilde{N}_{2}\in\mathbb{Z}_{+}$ such that for all $n\geq\widetilde{N}_{2}$, $i\in S_{n}$, $\mu\in\mathcal{M}(A)$,
$\text{(I) }+\text{ (III) }\leq\frac{\varepsilon}{4}+\dfrac{\varepsilon}{4(2M_{\varepsilon}+1)}.$
Below we estimate part (II). By the definition of the Prohorov metric $\rho$, we have that for any $\mu\in\mathcal{M}(A)$, $i\in I_{n}$,
$\rho\bigl{(}s(g^{n}),s(\mu,g^{n}_{-i})\bigr{)}\leq\sup\limits_{j\in I_{n}}\lambda_{n}(j).$
Since $\sup_{j\in I_{n}}\lambda_{n}(j)\rightarrow 0$, there exists $\widetilde{N}_{3}\in\mathbb{Z}_{+}$ such that for any $n\geq\widetilde{N}_{3}$, $\sup_{j\in I_{n}}\lambda_{n}(j)<\eta$. Thus, for any $n\geq\widetilde{N}_{3}$, $i\in I_{n}$, $\mu\in\mathcal{M}(A)$, $\rho\bigl{(}s(g^{n}),s(\mu,g^{n}_{-i})\bigr{)}<\eta.$ Recall that for all $i\in S_{n}$, $u_{i}^{n}$ is uniformly bounded by $M_{\varepsilon}$ and equicontinuous. As we proved in part (ii), we have that
[…]
for all $n\geq\widetilde{N}_{3}$, $i\in S_{n}$, $\mu\in\mathcal{M}(A)$. Thus for all $n\geq\max\{\widetilde{N}_{2},\widetilde{N}_{3}\}$, $i\in S_{n}$, $\mu\in\mathcal{M}(A)$,
[…]
@@
Step 5. For any $n\geq\max\{N_{\varepsilon},\widetilde{N}_{1},\widetilde{N}_{2},\widetilde{N}_{3}\}$, $i\in S_{n}$, $\mu\in\mathcal{M}(A)$, we have
$\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x^{n}_{i},\sum\limits_{j\in I_{n}}\lambda_{n}(j)\delta_{x^{n}_{j}}\bigr{)}\Bigr{]}$
@@
[…]
The first inequality follows from step 3 and the last inequality follows from step 4. The second inequality is due to the fact that $\widetilde{g}$ is a Nash equilibrium of $G$ such that for any player $i\in I$, her strategy $\widetilde{g}(i)$ is a best response with respect to the society summary s(g~)s($\widetilde${g}), and $G_{n}(i)\in G(I)$ for all $i\in I_{n}$, $n\in\mathbb{Z}_{+}$. Hence we have that
$\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x^{n}_{i},\sum\limits_{j\in I_{n}}\lambda_{n}(j)\delta_{x^{n}_{j}}\bigr{)}\Bigr{]}\geq\mathbb{E}\Bigl{[}u_{i}^{n}\bigl{(}x_{\mu},\lambda_{n}(i)\delta_{x_{\mu}}+\sum\limits_{j\in I_{n}\backslash\{i\}}\lambda_{n}(j)\delta_{x^{n}_{j}}\bigr{)}\Bigr{]}-\varepsilon,$
for all $n\geq\max\{N_{\varepsilon},\widetilde{N}_{1},\widetilde{N}_{2},\widetilde{N}_{3}\}$, $i\in S_{n}$, $\mu\in\mathcal{M}(A)$. Thus we conclude that $g^{n}$ is an $\varepsilon$-Nash equilibrium of $G_{n}$, for all $n\geq\max\{N_{\varepsilon},\widetilde{N}_{1},\widetilde{N}_{2},\widetilde{N}_{3}\}$.
Given any sequence $\{\varepsilon_{n}\}_{n\in\mathbb{Z}_{+}}$ such that $\varepsilon_{n}>0$ for all $n\in\mathbb{Z}_{+}$ and $\{\varepsilon_{n}\}_{n\in\mathbb{Z}_{+}}$ converges to 0, there exists a strictly increasing sequence $\{\overline{N}_{n}\}_{n\in\mathbb{Z}_{+}}$ such that $\overline{N}_{n}\in\mathbb{Z}_{+}$ for all $n\in\mathbb{Z}_{+}$, and $g^{m}$ is a $\varepsilon_{n}$-Nash equilibrium of $G_{m}$ for all $m\geq\overline{N}_{n}$. Let $\varepsilon_{k}=\varepsilon_{n}$ if $\overline{N}_{n}\leq k<\overline{N}_{n+1}$, for all $n,k\in\mathbb{Z}_{+}$. Thus we have that $\{\varepsilon_{k}\}_{k\in\mathbb{Z}_{+}}$ converges to 0, and $g^{k}$ is an $\varepsilon_{k}$-Nash equilibrium of $G_{k}$ for all $k\geq\overline{N}_{1}$, which completes our proof of Theorem 1.